OpenStudy (anonymous):

Translate to an equation: f $0.30 folders and n $1.50 notebooks total $12 Graph the equation and use the graph to determine three different combinations of folders and notebooks that total $12. Submit your graph through the dropbox.

OpenStudy (anonymous):

\[12=0.3f+1.5n\]

OpenStudy (anonymous):

All you need to do is pick three values for n, say, and obtain the corresponding f's from your equation/graph. For example, if you choose 2 notebooks, then n=2 and you have to find f:\[12=0.3f+1.5 \times 2 \rightarrow 9=0.3f \rightarrow f=9/0.3=30\]i.e. 30 folders can be bought with the remaining money. Hope it helps...

OpenStudy (anonymous):

For the example equation 12=0.3f+1.5 x 2, where does the 9 come from? I've been trying to do the calculations and I can't figure it out.